Let W = {(x, y, z): y² = x + z}, Is W a subspace of R³

1
Expert's answer
2022-04-07T12:08:59-0400

"W" consists of all vectors "(x,y,z)" satisfying: "y^2=x+z". Suppose that "(x_1,y_1,z_1)" and "(x_2,y\n_2,z_2)" belong to "W". It means that "y_1^2=x_1+z_1" and "y_2^2=x_2+z_2". But, it does not mean that "(x_1+x_2,y_1+y_2,z_1+z_2)" belongs to "W". I.e., it may not satisfy "(y_1+y_2)^2=(x_1+x_2)+(z_1+z_2)". For example, we can choose "(1,2,3)" and "(0,1,1)". It is clear that both vectors belong to "W". But "(1,3,4)" does not belong to "W", since "3^2\neq1+4". The latter means that one of subspace properties is not satisfied for "W" . Therefore, it is not a subspace of "\mathbb{R}^3" .